OpenStudy (anonymous):

Need help. What is X/X^2-1 subtract 2/X+1

OpenStudy (campbell_st):

start by factorising the denominator of the 1st fraction its the difference of 2 squares. \[\frac{x}{(x -1)(x + 1)} - \frac{2}{x + 1}\] you will need a common denominator so that you can subtract the fractions so the 2nd fraction needs to be multiplied by (x-1)/(x -1) \[\frac{x}{(x -1)(x +1)} - \frac{2}{(x +1)} \times \frac{(x -1)}{(x - 1)}\] this will result in \[\frac{x}{(x -1)(x+1)} - \frac{2(x -1)}{(x -1)(x +1)}\] since the denominators are the same, just work on simplifying the numerators.